OpenStudy (anonymous):

a ladder is 6m long. if the height of the top of the ladder must be no greater than 10 times the distance from the base to the wall, how high up a wall can the top of the ladder be placed.

OpenStudy (anonymous):

x = dist ladder from the building 10x = height of the ladder on the bldg x^2 + (10x)^2 = 6^2 x^2 + 100x^2 = 36 101x^2 = 36 x^2 = 36/101 x =√36/101 x=.597 meters 10*.597 = 5.97 meters